# Giant Molecular clouds in RCW 106 (G333): Galactic mini-starbursts and massive star formation induced by supersonic cloud-cloud collisions

### v1

@@ Abstract
To reveal the origin of the mini-starbursts in the Milky Way, we carried out large-scale CO observations toward the RCW 106 giant molecular cloud (GMC) complex using the NANTEN2 4-m radio telescope operated by Nagoya University. We also analyzed the Mopra Southern Galactic plane CO survey and Herschel infrared continuum archival data. The RCW 106 GMC complex contains the radial velocity components of $-68$ km s-1 and $-50$ km s-1 reported by Nguyen et al. (2015). Focusing on the RCW 106 East and West region with the massive star formation having the bright infrared dust emission, we found that these regions have three different velocity components with $\sim$ 10 km s-1 differences. The two out of three velocity components show morphological correspondence with the infrared cold dust emission and connect with the bridge feature on a position-velocity diagram. Therefore, two molecular clouds (MCs) with $\sim$ 10 km s-1 differences are likely to be physically associated with massive star-forming regions in the GMC complex. Based on these observational results, we argue that mini-starbursts and massive star/cluster formation in the RCW 106 GMC complex are induced by supersonic cloud-cloud collisions in an agglomerate of molecular gas on the Scutum-Centurus arm.
@@ 3.1 CO spatial and velocity distributions of the RCW 106 GMC complex
Figures 3 (a) and (b) show the 12CO $J$ = 1–0 spectra in RCW 106 East and RCW 106 West, respectively. The positions of the spectra are shown as the orange cross marks in Figures 5a and 5b. Previous studies indicate that the velocity range from $-110$ km s-1 to $-90$ km s-1 is associated with the Norma arm with a distance of $\sim$ 4.6–7.5 kpc and the background component toward RCW 106 (Merello et al., 2013; Nguyen et al., 2015). The velocity range from $-80$ km s-1 to $-30$ km s-1 is associated with RCW 106 located on the Scutum-Centaurus arm and is composed of two velocity components, which are from $-75$ km s-1 to $-60$ km s-1 and from $-60$ km s-1 to $-40$ km s-1 . Hereafter, we call these components “$-68$ km s-1 GMC” and “$-50$ km s-1 GMC”.
@@
Figure 4: The 12CO $J$ = 1–0 velocity chennel map obtained by NANTEN2. The cube data is smoothed to the velocity domain of 9 channels by the hanning window. Red dotted rectangles show giant molecular filaments of GMF 335.6−333.6a and GMF 335.6−333.6b.
@@ 4.1 The scenario of massive stars/cluster formation in the RCW 106 GMC complex
Our detailed analysis of RCW 106 East and RCW 106 West indicates that the two MCs having a velocity difference of $\sim 10$ km s-1 are physically associated with massive star-forming regions in the RCW 106 GMC complex. Figures 11(a) and 11(b) show the overlaid spatial distributions of two MCs focusing on RCW 106 East and RCW 106 West, respectively. These MCs overlap around massive star-forming regions of G333.6-0.2, IRAS 16172-5028, G333.0-0.4 in RCW 106 East, and IRAS 16065-5158 in RCW 106 West. This observational feature is consistent with the signature of cloud-cloud collisions reported by previous studies of Galactic min-starburst region NGC 6334 (see Figure 8 in Fukui et al., 2018b). Figure 12 (a) and (b) show the position-velocity diagrams focusing on the massive star-forming regions in RCW 106 East and West, respectively. We can find the bridge features connecting two CO velocity components on the velocity space. If two molecular clouds collide and interact, we can observe the bridging feature connecting in the velocity space on the position-velocity diagram by comparing CO observations with synthetic observations (Haworth et al., 2015a, b; Torii et al., 2017; Bisbas et al., 2017). We suggest that these results indicate the signature of the interaction between two molecular clouds in RCW 106 East and West. We comment on the viewing angle effects and expected velocity of the star-forming regions formed by colliding two clouds. Fukui et al. (2018a) argued the projection and viewing angle effects distort the contact interface layer on the position-velocity diagram comparing cloud-cloud collision in the Orion nebula cluster with the simulation model by Takahira et al. (2014) and Haworth et al. (2015a, b). In RCW106 East, the Higal compact sources have velocities around $-53$  km s-1 , and in RCW106 West, they have velocities around $-65$  km s-1 (Mège et al., 2021). In the cloud-cloud collision model with internal turbulent motion, different sizes, and densities, the shock-compressed layer formed by momentum exchange does not necessarily have to be an intermediate velocity of the two colliding clouds (e.g., Takahira et al., 2014; Haworth et al., 2015b). In addition, the velocities of massive clumps formed by cloud-cloud collision can have offsets from the intermediate velocity of two colliding clouds because shock compression enhances turbulence in the compressed layer (Inoue & Fukui, 2013; Fukui et al., 2021b).
@@
Figure 13 shows the schematic picture of our proposed scenario. Figure 13(a) illustrates the location of the RCW 106 GMC complex in the Scutum-Centarus arm of the Milky Way. Molecular clouds, whose sizes and masses of $\sim$ 20–30 pc and $\sim 10^{4}$$10^{6}M_{\odot}$ collide with the supersonic velocity of $\sim 10$ km s-1 , as shown in Figure 13(b). The shock compression by supersonic collisions of molecular clouds produces dense cores and triggers massive star/cluster formation (see the review paper of Fukui et al., 2021a). Indeed, the numerical simulations show that supersonic cloud-cloud collisions having different sizes and densities can produce massive molecular cloud cores required for massive star formation at the interface layer by shock compression (Habe & Ohta, 1992; Anathpindika, 2010; Takahira et al., 2014, 2018; Shima et al., 2018). Recent simulations considering magnetic fields also reveal massive dense cores produced by cloud-cloud collisions (Inoue & Fukui, 2013; Wu et al., 2015; Fukui et al., 2021b; Sakre et al., 2021; Kinoshita & Nakamura, 2022; Kashiwagi et al., 2023).
@@
Figure 13: (a) The location at the RCW 106 GMC complex superposed on the face-on illustration of the Milky Way (NASA/JPL-Caltech/ESO/R. Hurt). The distance to the Galactic center ($R_{0}$) is adapted from the Very Long Baseline Interferometry (VLBI) astrometry results (Reid et al., 2019; VERA Collaboration et al., 2020). (b) Schematic illustration of our proposed massive star/cluster formation scenario in the RCW 106 GMC complex.
@@ 4.2 Cloud-cloud collisions in the RCW 106 GMC complex
4.3 Comparision with properties of Galactic mini-starburst regions associated with the GMC complex.
Table 3: Comparision with the Galactic mini-starbursts induced by cloud-cloud collisions in the GMC complex.
@@ 5 Conclusion
    The two out of three velocity components in RCW 106 East and RCW 106 West coincide CO peaks with infrared dust emission and connect each other with the bridge feature on the position-velocity diagram. Thus, these two velocity components having a velocity difference of $\sim$ 10 km s-1  are likely to be physically associated with massive star-forming regions and interact with each other in the RCW 106 GMC complex. We suggest that massive star/cluster formation in the RCW 106 GMC complex might be induced by supersonic cloud collisions in the Scutum-Centurus arm.
@@
    The RCW 106 East has the peak H2 column density of $N({\rm H_{2}})>10^{23}$ cm-2. This column density is satisfied with the initial condition of $>10$ O-type star formation reported by previous studies of other massive star-forming regions (Enokiya et al., 2021). The total mass of RCW 106 East is $\sim 2\times 10^{6}\ M_{\odot}$, which is massive compared with other mini-starburst regions in the Milky Way. We point out that cloud-cloud collisions frequently occur in the RCW 106 GMC complex on the Scutum-Centarus arm, which might trigger extreme star/cluster formation.

### v2

@@ Abstract
To reveal the origin of the mini-starbursts in the Milky Way, we carried out large-scale CO observations toward the RCW 106 giant molecular cloud (GMC) complex using the NANTEN2 4-m radio telescope operated by Nagoya University. We also analyzed the Mopra Southern Galactic plane CO survey and Herschel infrared continuum archival data. The RCW 106 GMC complex contains the radial velocity components of $-68$ km s-1 and $-50$ km s-1 reported by Nguyen et al. (2015). Focusing on the RCW 106 East and West region with the massive star formation having the bright infrared dust emission, we found that these regions have three different velocity components with $\sim$ 10 km s-1 differences. The two out of three velocity components show morphological correspondence with the infrared cold dust emission and connect with the bridge feature on a position-velocity diagram. Therefore, two molecular clouds (MCs) with $\sim$ 10 km s-1 differences are likely to be physically associated with massive star-forming regions in the GMC complex. Based on these observational results, we argue that mini-starbursts and massive star/cluster formation in the RCW 106 GMC complex are induced by supersonic cloud-cloud collisions in an agglomerate of molecular gas on the Scutum-Centaurus arm.
@@ 3.1 CO spatial and velocity distributions of the RCW 106 GMC complex
Figures 3 (a) and (b) show the 12CO $J$ = 1–0 spectra in RCW 106 East and RCW 106 West, respectively. The positions of the spectra are shown as the orange cross marks in Figures 5a and 5b. Previous studies indicate that the velocity range from $-110$ km s-1 to $-90$ km s-1 is associated with the Norma arm with a distance of $\sim$ 4.6–7.5 kpc and the background component toward RCW 106 (Merello et al., 2013; Nguyen et al., 2015). The velocity range from $-80$ km s-1 to $-30$ km s-1 is associated with RCW 106 located on the Scutum-Centaurus arm and is composed of two velocity components, which are from $-75$ km s-1 to $-60$ km s-1 and from $-60$ km s-1 to $-45$ km s-1 . Hereafter, we call these components “$-68$ km s-1 GMC” and “$-50$ km s-1 GMC”.
@@
Figure 4: The 12CO $J$ = 1–0 velocity channel map obtained by NANTEN2. The cube data is smoothed to the velocity domain of 9 channels by the hanning window. Red dotted rectangles show giant molecular filaments of GMF 335.6−333.6a and GMF 335.6−333.6b.
@@ 4.1 The scenario of massive stars/cluster formation in the RCW 106 GMC complex
Our detailed analysis of RCW 106 East and RCW 106 West indicates that the two MCs having a velocity difference of $\sim 10$ km s-1 are physically associated with massive star-forming regions in the RCW 106 GMC complex. Figures 11(a) and 11(b) show the overlaid spatial distributions of two MCs focusing on RCW 106 East and RCW 106 West, respectively. These MCs overlap around massive star-forming regions of G333.6-0.2, IRAS 16172-5028, G333.0-0.4 in RCW 106 East, and IRAS 16065-5158 in RCW 106 West. This observational feature is consistent with the signature of cloud-cloud collisions reported by previous studies of Galactic min-starburst region NGC 6334 (see Figure 8 in Fukui et al., 2018b). Figures 12 (a) and (b) show the position-velocity diagrams focusing on the massive star-forming regions in RCW 106 East and West, respectively. We can find the bridge features connecting two CO velocity components on the velocity space. If two molecular clouds collide and interact, we can observe the bridging feature connecting in the velocity space on the position-velocity diagram by comparing CO observations with synthetic observations (Haworth et al., 2015a, b; Torii et al., 2017; Bisbas et al., 2017). We suggest that these results indicate the signature of the interaction between two molecular clouds in RCW 106 East and West. We comment on the viewing angle effects and expected velocity of the star-forming regions formed by colliding two clouds. Fukui et al. (2018a) argued the projection and viewing angle effects distort the contact interface layer on the position-velocity diagram comparing cloud-cloud collision in the Orion nebula cluster with the simulation model by Takahira et al. (2014) and Haworth et al. (2015a, b). In RCW106 East, the Higal compact sources have velocities around $-53$  km s-1 , and in RCW106 West, they have velocities around $-65$  km s-1 (Mège et al., 2021). In the cloud-cloud collision model with internal turbulent motion, different sizes, and densities, the shock-compressed layer formed by momentum exchange does not necessarily have to be an intermediate velocity of the two colliding clouds (e.g., Takahira et al., 2014; Haworth et al., 2015b). In addition, the velocities of massive clumps formed by cloud-cloud collision can have offsets from the intermediate velocity of two colliding clouds because shock compression enhances turbulence in the compressed layer (Inoue & Fukui, 2013; Fukui et al., 2021b).
@@
Figure 13 shows the schematic picture of our proposed scenario. Figure 13(a) illustrates the location of the RCW 106 GMC complex in the Scutum-Centaurus arm of the Milky Way. Molecular clouds, whose sizes and masses of $\sim$ 20–30 pc and $\sim 10^{4}$$10^{6}M_{\odot}$ collide with the supersonic velocity of $\sim 10$ km s-1 , as shown in Figure 13(b). The shock compression by supersonic collisions of molecular clouds produces dense cores and triggers massive star/cluster formation (see the review paper of Fukui et al., 2021a). Indeed, the numerical simulations show that supersonic cloud-cloud collisions having different sizes and densities can produce massive molecular cloud cores required for massive star formation at the interface layer by shock compression (Habe & Ohta, 1992; Anathpindika, 2010; Takahira et al., 2014, 2018; Shima et al., 2018). Recent simulations considering magnetic fields also reveal massive dense cores produced by cloud-cloud collisions (Inoue & Fukui, 2013; Wu et al., 2015; Fukui et al., 2021b; Sakre et al., 2021; Kinoshita & Nakamura, 2022; Kashiwagi et al., 2023).
@@
Figure 13: (a) The location at the RCW 106 GMC complex superposed on the face-on illustration of the Milky Way (NASA/JPL-Caltech/ESO/R. Hurt). The distance to the Galactic center ($R_{0}$) is adopted from the Very Long Baseline Interferometry (VLBI) astrometry results (Reid et al., 2019; VERA Collaboration et al., 2020). (b) Schematic illustration of our proposed massive star/cluster formation scenario in the RCW 106 GMC complex.
@@ 4.2 Cloud-cloud collisions in the RCW 106 GMC complex
4.3 Comparison with properties of Galactic mini-starburst regions associated with the GMC complex.
Table 3: Comparison with the Galactic mini-starbursts induced by cloud-cloud collisions in the GMC complex.
@@ 5 Conclusion
    The two out of three velocity components in RCW 106 East and RCW 106 West coincide CO peaks with infrared dust emission and connect each other with the bridge feature on the position-velocity diagram. Thus, these two velocity components having a velocity difference of $\sim$ 10 km s-1  are likely to be physically associated with massive star-forming regions and interact with each other in the RCW 106 GMC complex. We suggest that massive star/cluster formation in the RCW 106 GMC complex might be induced by supersonic cloud collisions in the Scutum-Centaurus arm.
@@
    The RCW 106 East has the peak H2 column density of $N({\rm H_{2}})>10^{23}$ cm-2. This column density is satisfied with the initial condition of $>10$ O-type star formation reported by previous studies of other massive star-forming regions (Enokiya et al., 2021). The total mass of RCW 106 East is $\sim 2\times 10^{6}\ M_{\odot}$, which is massive compared with other mini-starburst regions in the Milky Way. We point out that cloud-cloud collisions frequently occur in the RCW 106 GMC complex on the Scutum-Centaurus arm, which might trigger extreme star/cluster formation.